# A Survey on Self-play Methods in Reinforcement Learning

### v3

@@ 2.2.3 Transitive Game and Non-transitive Game
For the sake of simplicity, we restrict our focus to two-player zero-sum symmetric games. In a transitive game, the strategies or outcomes follow a transitive relationship. Formally, $\forall\pi_{i},\pi_{j},\pi_{k}\in\Pi$, if $u(\pi_{i},\pi_{j})>0$ and $u(\pi_{j},\pi_{k})>0$, then it must follow that $u(\pi_{i},\pi_{k})>0$. This transitive property simplifies the strategic landscape, allowing for an ordinal ranking of strategies. Conversely, in a non-transitive game, $\exists~{}\pi_{i},\pi_{j},\pi_{k}\in\Pi$ such that $u(\pi_{i},\pi_{j})>0$ and $u(\pi_{j},\pi_{k})>0$, but $u(\pi_{i},\pi_{k})\leq 0$. This introduces a cyclic relationship among strategies, thereby complicating the game. The complexity often results in a mixed-strategy equilibrium, where players randomize their choices among multiple strategies to maximize their expected payoff. A classical example of a non-transitive game is Rock-Paper-Scissors, in which no single strategy uniformly dominates all others.
2.2.4 Stage Game and Repeated Game
@@ 2.2.5 Nash Equilibrium
$BR_{i}(\pi_{-i})=\arg\max_{\pi_{i}}u_{i}(\pi_{i},\pi_{-i}).$ (1)
A strategy $\pi_{i}^{*}$ is an $\bm{\epsilon}$-BR to strategies $\pi_{-i}$ if:
$u_{i}(\pi_{i}^{*},\pi_{-i})\geq u_{i}(BR_{i}(\pi_{-i}),\pi_{-i})-\epsilon,$ (2)
@@
A strategy profile $(\pi_{1}^{*},\pi_{2}^{*},...,\pi_{n}^{*})$ is an $\bm{\epsilon}$-NE if, for each player $i$:
$u_{i}(\pi_{i}^{*},\pi_{-i}^{*})\geq u_{i}(\pi_{i},\pi_{-i}^{*})-\epsilon,\forall\pi_{i},$ (4)
@@
However, computing NE is generally intractable in complex games, so some researchers utilize $\alpha$-Rank (Omidshafiei et al., 2019) and Correlated Equilibrium (CE) (Aumann, 1974) as alternatives. Some studies also resort to Replicator Dynamics (Taylor & Jonker, 1978) to analyze strategy evolution.
3 Algorithms
@@
Based on existing self-play work (Hernandez et al., 2021; Lanctot et al., 2017; Liu et al., 2022c; Garnelo et al., 2021), we propose a self-play framework (Algo. 1) that boasts enhanced expressivity and superior generalization capabilities. For simplicity, our framework is illustrated for symmetric games. All players share a policy population with a fixed maximum size. In each iteration, a newly initialized policy is trained, while opponent policies are sampled from the population. After training, the new policy is added to the population. The updated policy population is then evaluated, and then the opponent sampling strategy is recalculated for the next iteration.
@@
This section is organized as follows: In Sec.3.1, we provide a formalized description of our framework. We then categorize self-play algorithms into four primary groups: traditional self-play algorithms (Sec. 3.2), the PSRO series (Sec. 3.3), the ongoing-training-based series (Sec. 3.4), and the regret-minimization-based series (Sec. 3.5). We analyze how these four categories align with our framework and introduce corresponding algorithms for each category. To make it more straightforward, we highlight the representative classic algorithms from these and present them in Table 1 for comparison within our framework. In Sec. 3.6, we compare the four categories, explain our classification rationale, and highlight the distinctions between our framework and existing ones to demonstrate its greater generality.
3.1 Framework Definition
@@
Figure 3: Illustration of our framework. The top row shows two different initialization methods for the policy population $\Pi$ and the interaction matrix $\Sigma$: lazy initialization and immediate initialization (corresponding to Line 1 in Algo. 1). The bottom row illustrates two training paradigms: standard training and ongoing training (corresponding to Lines 2~9 in Algo. 1). We categorize self-play algorithms into four types: traditional self-play, the PSRO series, the ongoing-training-based series, and the regret-minimization-based series. Among these, traditional self-play, the PSRO series, and the regret-minimization-based series utilize lazy initialization and standard training, whereas the ongoing-training-based series employs immediate initialization and ongoing training. For a detailed description and analysis, please refer to Sec.3 and Table1.
@@
3:     for $\sigma_{i}\in\Sigma$ do
4:        Initialize $\pi_{i}^{h}$
5:        $\pi_{i}^{h}\leftarrow\textit{ORACLE}(\pi_{i}^{h},\sigma_{i},\Pi)$$\triangleright$ Compute the oracle.
6:        $\mathcal{P}\leftarrow\textit{EVAL}(\Pi)$triangleright Get policies’ performance.
7:        $\Sigma\leftarrow\textit{MSS}(\mathcal{P})$$\triangleright$ Update the interaction matrix.
@@
  • $\Pi:=\{\pi_{i}(\cdot|h(i))\}_{i=1}^{N}$: Each policy $\pi_{i}$ in the policy population $\Pi$ is conditioned on a policy condition function $h(i)$, which provides supplementary information for certain algorithms. We denote the policy $\pi_{i}(\cdot|h(i))$ as $\pi_{i}^{h}$. Note that $i$ refers to the $i$th policy in the population, not the $i$th player, and $N$ denotes the policy population size, not the number of players. $\Pi$ can be initialized (Line 1 in Algo. 1) in two ways: lazy initialization and immediate initialization. In lazy initialization, $\Pi$ starts with $N$ placeholder policies, with the actual policies being initialized during the training iterations (Line 4 in Algo. 1). In immediate initialization, $\Pi$ is initialized upfront with $N$ real policies, which may be randomly generated or pre-trained models. A policy $\pi_{i}^{h}$ is considered inactive if it is a placeholder; otherwise, it is considered active.
@@
  • $\Sigma:=\{\sigma_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{1}}$: The interaction matrix $\Sigma$ consists of rows $\sigma_{i}$, where each $\sigma_{i}\in\mathbb{R}^{C_{1}}$ represents the opponent sampling strategy of policy $i$, indicating how to sample opponents’ policies. The dimension $C_{1}$ varies according to the sampling method. For instance, $\sigma_{i}$ can represent the probability distribution over opponent policies in the policy population $\Pi$, where $C_{1}=N^{n-1}$ and $n$ denotes the number of players. Alternatively, $\sigma_{i}$ can be viewed as parameters of a sampling network. Specially, in a two-player game, if $C_{1}=N$ and $\sigma_{ij}$ represents the probability that policy $i$ is optimized against policy $j$, $\Sigma$ can be depicted in directed interaction graphs (Fig. 4). $\Sigma$ will be initialized (Line 1 in Algo. 1) based on the initialization of $\Pi$. If $\Pi$ uses lazy initialization, $\Sigma$ will be initialized as an empty matrix. If $\Pi$ uses immediate initialization, $\Sigma$ will be calculated based on $\Pi$. We will discuss this in more detail later.
@@
  • For epoch $e\in[[E]]$ (Line 2 in Algo. 1): $E$ denotes the total number of epochs for the entire policy population. For example, if the algorithm only introduces new policies into the population without updating existing ones, then $E=1$. This implies that only the inactive policy is likely to be chosen for training in each iteration and turned into an active one, while activa policies remain unchanged. Conversely, if active policies are updated multiple times throughout the algorithm, then $E>1$. Indeed, $E$ accurately reflects the number of updates performed.
@@
  • Initialize $\pi_{i}^{h}$ (Line 4 in Algo. 1): The initialization of $\pi_{i}^{h}$ can vary depending on the algorithm being used. It can be initialized randomly or by leveraging pre-trained models (Silver et al., 2016; Vinyals et al., 2019) or through a recently updated policy.
@@
  • $\text{ORACLE}(\pi_{i},\sigma_{i},\Pi)$ (Line 5 in Algo. 1): The ORACLE is an abstract computational entity that returns a new policy adhering to specific criteria. Here, we divide ORACLE into three types. (1) One type is the BR oracle, which is designed to identify the optimal counter-strategies against an opponent’s strategy, including finding NE (McMahan et al., 2003). However, it often requires considerable computational effort. (2) To alleviate the computational demands, the approximate best response (ABR) oracle is introduced, using techniques such as RL (Algo. 2), evolution-theory-based methods (Algo. 3) or regret minimization methods (Algo. 4). (3) Other specially crafted ORACLEs are either tailored to specific meta-strategy solvers (MSSes) (Muller et al., 2020; Marris et al., 2021) or introduced to enhance diversity (Balduzzi et al., 2019; Perez-Nieves et al., 2021; Liu et al., 2021). Some details of Algo. 23 and 4 need to be mentioned. Algo. 2 is a RL method that can be widely used across different categories. Moreover, off-policy RL algorithms typically require a replay buffer to gather samples; however, they are not explicitly included for simplicity. Algo. 3 is specifically utilized by Regularized Nash Dynamics (R-NaD) (Perolat et al., 2021) and Algo. 4 is specifically utilized by the regret-minimization-based series. They will be introduced in detail in Sec. 3.3.5 and Sec. 3.5 respectively.
@@
  • $\text{EVAL}(\Pi)$ (Line 6 in Algo. 1): Evaluating the policy population $\Pi$. There are multiple evaluation metrics available to assess the performance of each policy. The performance matrix is represented as $\mathcal{P}:=\{p_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times C_{2}}$. $p_{i}$ is the performance of policy i, and the dimension $C_{2}$ depends on the evaluation metric used. For instance, $p_{i}$ can be depicted as the relative skill like Elo ratings ($C_{2}=1$) or can be depicted as the payoff tensor ($C_{2}=N^{n-1}$, where n is the number of players). Specially, in two-player symmetric zero-sum games, the expected payoffs can serve as the evaluation metric. In such cases, $\mathcal{P}$ is a square matrix ($C_{2}=N$).
@@
  • $\textit{MSS}:\mathbb{R}^{N\times C_{2}}\rightarrow\mathbb{R}^{N\times C_{1}}$ (Line 7 in Algo. 1): A meta-strategy solver (MSS) MSS takes the performance matrix $\mathcal{P}$ as its input and produces a new interaction matrix $\Sigma$ as its output. If the MSS produces a constant matrix irrespective of the input, the evaluation step (Line 6 in Algo. 1) can be skipped to reduce computations. Moreover, if $\Pi$ uses immediate initialization, $\Sigma$ will be initialized (Line 1 in Algo. 1) by first evaluating the performance and then applying the MSS.
@@
Figure 4: Interaction matrix $\Sigma$ examples. When $C_{1}=N$ and the opponent sampling strategy $\sigma_{ij}$ represents the probability that policy $i$ is optimized against policy $j$, the interaction matrix $\Sigma$ can be depicted in directed interaction graphs. Here, we consider three representative self-play algorithms. In the Top section, we define the interaction matrix $\Sigma\in\mathbb{R}^{3\times 3}$ as $\{\sigma_{i}\}_{i=1}^{3}$. In the Bottom section, we present directed interaction graphs where the outgoing edges from each node are equally weighted, and their weights collectively sum to one. The relationship between the Top and Bottom sections is established through directed edges: an edge directed from node $i$ to node $j$ with a weight of $\sigma_{ij}$ signifies that policy $i$ is optimized against policy $j$ with a probability of $\sigma_{ij}$. Note that this figure is reproduced from Liu et al. (2022c) and this concept is initially proposed by Garnelo et al. (2021).
@@
0:  $\pi_{i}^{h}$$\triangleright$ Policy $i$ is being trained.
0:  $\sigma_{i}$$\triangleright$ Opponent sampling strategy of policy $i$.
0:  $\Pi$$\triangleright$ Policy population.
1:  while $\pi_{i}^{h}$ is not valid do
@@
3:        sample $\bm{\pi_{opp}^{h}}\sim P(\sigma_{i})$$\triangleright$ Policies of opponents.
4:        $\bm{\pi}=(\pi_{i}^{h},\bm{\pi_{opp}^{h}})$
5:        $s_{0},\bm{o_{0}}\sim\rho$triangleright Initial state.
@@
7:           $\bm{a_{t}}\sim\bm{\pi}(\bm{o_{t}})$
8:           $s_{t+1},\bm{o_{t+1}}\sim P(s_{t},\bm{a_{t}})$
9:           $\bm{r_{t}}\leftarrow\bm{R}(s_{t},\bm{a_{t}})$
@@
11:        $\pi_{i}^{h}\leftarrow\text{update}(\pi_{i}^{h})$$\triangleright$ Using RL algorithms
12:     end for
13:  end while
14:  return  $\pi_{i}^{h}$
@@
0:  $\pi_{i}^{h}$$\triangleright$ Policy $i$ is being trained.
0:  $\sigma_{i}$$\triangleright$ Opponent sampling strategy of policy $i$.
0:  $\Pi$$\triangleright$ Policy population.
1:  sample $\bm{\pi_{opp}^{h}}\sim P(\sigma_{i})$$\triangleright$ Policies of opponents.
2:  $\bm{\pi_{\text{reg}}}=(\pi_{i}^{h},\bm{\pi_{opp}^{h}})$$\triangleright$ Regularization policies.
3:  Transformed the reward according to $\bm{\pi_{\text{reg}}}$.
4:  $\pi_{i}^{h}\leftarrow$ Use replicator dynamics to play the reward-transformed game until convergence.
5:  return  $\pi_{i}^{h}$
@@
0:  $\pi_{i}^{h}$$\triangleright$ Policy $i$ is being trained.
0:  $\sigma_{i}$$\triangleright$ Opponent sampling strategy of policy $i$.
0:  $\Pi$$\triangleright$ Policy population.
1:  for each player j do
2:     sample $\bm{\pi_{opp}^{h}}\sim P(\sigma_{i})$$\triangleright$ Policies of opponents.
3:     $\bm{\pi}=(\pi_{i}^{h}(j),\bm{\pi_{opp}^{h}}(-j))$
4:     Use regret matching to play the game and update $h(i)$ with new regret minimization information.
@@
6:  return  $\pi_{i}^{h}$
3.2 Traditional Self-play Algorithms

### v4

@@ 2.2.2. Transitive Game and Non-Transitive Game
To provide a clearer introduction to transitive and non-transitive games, in this section, we confine our analysis to two-player zero-sum symmetric games. In a transitive game, the strategies or outcomes follow a transitive relationship. Formally, $\forall\pi_{i},\pi_{j},\pi_{k}\in\Pi$, if $u(\pi_{i},\pi_{j})>0$ and $u(\pi_{j},\pi_{k})>0$, then it must follow that $u(\pi_{i},\pi_{k})>0$. This transitive property simplifies the strategic landscape, allowing for an ordinal ranking of strategies. Conversely, in a non-transitive game, $\exists~\pi_{i},\pi_{j},\pi_{k}\in\Pi$ such that $u(\pi_{i},\pi_{j})>0$ and $u(\pi_{j},\pi_{k})>0$, but $u(\pi_{i},\pi_{k})\leq 0$. This introduces a cyclic relationship among strategies, thereby complicating the game. The complexity often results in a mixed-strategy equilibrium, where players randomize their choices among multiple strategies to maximize their expected payoff. A classical example of a non-transitive game is Rock-Paper-Scissors, in which no single strategy uniformly dominates all others.
2.2.3. Stage Game and Repeated Game
@@ 2.2.4. Nash Equilibrium
(1) $BR_{i}(\pi_{-i})=\arg\max_{\pi_{i}}u_{i}(\pi_{i},\pi_{-i}).$
A strategy $\pi_{i}^{*}$ is an $\boldsymbol{\epsilon}$-BR to strategies $\pi_{-i}$ if:
(2) $u_{i}(\pi_{i}^{*},\pi_{-i})\geq u_{i}(BR_{i}(\pi_{-i}),\pi_{-i})-\epsilon,$
@@
A strategy profile $(\pi_{1}^{*},\pi_{2}^{*},...,\pi_{n}^{*})$ is an $\boldsymbol{\epsilon}$-NE if, for each player $i$:
(4) $u_{i}(\pi_{i}^{*},\pi_{-i}^{*})\geq u_{i}(\pi_{i},\pi_{-i}^{*})-\epsilon,\forall\pi_{i},$
@@
However, computing NE is generally intractable in complex games, so some researchers utilize $\alpha$-Rank (Omidshafiei et al., 2019) and Correlated Equilibrium (CE) (Aumann, 1974) as alternatives. Some studies also resort to Replicator Dynamics (Taylor and Jonker, 1978) to analyze strategy evolution.
2.2.5. Static Game and Dynamic Game
The market entry game is a classic example in game theory. There are two players: Firm 1 (the entrant), which is considering whether to enter a market, and Firm 2 (the incumbent), which currently dominates that market. The game outcomes are as follows:
  • If the entrant stays out of the market, the entrant earns a payoff of 4, and the incumbent earns 10.
  • If the entrant enters and the incumbent chooses to attack, both firms receive a payoff of 3.
  • If the entrant enters and the incumbent chooses not to attack, both firms receive a payoff of 6.
Figure 2. The example of the market entry game. (a) Matrix representation of simultaneous market entry game in normal-form. (b) Game tree representation of sequential market entry game in extensive-form.
In the simultaneous version of the market entry game, both players choose their strategies at the same time. The entrant decides whether to enter, and the incumbent simultaneously decides whether to attack, without knowing the entrant’s decision. Games of this kind are referred to as static games. Although a static game can be represented in either normal- or extensive-form, it is conventionally shown in the normal form because it succinctly captures the strategic structure of simultaneous moves (as shown in Fig. 2a). In the sequential version, the entrant moves first, and the incumbent responds after observing the entrant’s decision. Games of this nature are called dynamic games. Although any dynamic game admits a normal-form representation, that representation scales exponentially with the number of information sets. Consequently, dynamic games are typically illustrated in extensive form, whose game tree compactly encodes the order of moves and the information available to each player (as depicted in Fig. 2b).
The market entry game demonstrates how the structure of a game influences equilibrium outcomes. In the static version, non-credible threats—such as the incumbent’s threat to attack even when it is not in its best interest—may support a Nash equilibrium where the entrant stays out. In contrast, the dynamic version allows for refinement through subgame perfect Nash equilibrium (SPNE), which requires that players’ strategies constitute a Nash equilibrium in every subgame. Under this refinement, the threat to attack is ruled out as non-credible, and the predicted outcome is that the entrant will enter and the incumbent will accommodate, since attacking would harm the incumbent itself.
3. Algorithms
@@
Based on existing self-play work (Hernandez et al., 2021; Lanctot et al., 2017; Liu et al., 2022c; Garnelo et al., 2021), we propose a self-play framework (Algo. 1) that boasts enhanced expressivity and superior generalization capabilities. For simplicity, our framework is illustrated for symmetric games. All players share a policy population with a fixed maximum size. In each iteration, a newly initialized policy is trained, while opponent policies are sampled from the population. After training, the new policy is added to the population, possibly replacing an existing policy. The updated policy population is then evaluated, and then the opponent sampling strategy is recalculated for the next iteration.
@@
This section is organized as follows: In Sec.3.1, we provide a formalized description of our framework. We then categorize self-play algorithms into four primary groups: traditional self-play algorithms (Sec. 3.2), the PSRO series (Sec. 3.3), the ongoing-training-based series (Sec. 3.4), and the regret-minimization-based series (Sec. 3.5). We analyze how these four categories align with our framework and introduce corresponding algorithms for each category. To make it more straightforward, we highlight the representative classic algorithms from these and present them in Table 1 for comparison within our framework. Sec. 3.6 compares the four categories, explains our classification rationale, contrasts the framework with prior ones to demonstrate its greater generality, and highlights the design insights it offers for future self-play research.
3.1. Framework Definition
@@
Figure 3. Illustration of our framework. The top row shows two different initialization methods for the policy population $\Pi$ and the interaction matrix $\Sigma$: lazy initialization and immediate initialization (corresponding to Line 1 in Algo. 1). The bottom row illustrates two training paradigms: standard training and ongoing training (corresponding to Lines 2~9 in Algo. 1). We categorize self-play algorithms into four types: traditional self-play, the PSRO series, the ongoing-training-based series, and the regret-minimization-based series. Among these, traditional self-play, the PSRO series, and the regret-minimization-based series utilize lazy initialization and standard training, whereas the ongoing-training-based series employs immediate initialization and ongoing training. For a detailed description and analysis, please refer to Sec.3 and Table1.
@@
3:  for $\sigma_{[k]}\in[[K]]$ do
4:   Initialize $\pi_{{[k]}}^{h}$
5:   $\pi_{[k]}^{h}\leftarrow\textit{ORACLE}(\pi_{[k]}^{h},\sigma_{[k]},\Pi)\triangleright$ Compute the oracle.
6:   $\mathcal{P}\leftarrow\textit{EVAL}(\Pi)$ $\triangleright$ Get policies’ performance (optional).
7:   $\Sigma\leftarrow\textit{MSS}(\mathcal{P})\triangleright$ Update the interaction matrix (optional).
@@
  • $\Pi:=\{\pi_{[k]}(\cdot|h(k)\}_{k=1}^{K}$: Each policy $\pi_{[k]}$ in the policy population $\Pi$ is conditioned on a policy condition function $h(k)$, which provides supplementary information for certain algorithms. We denote the policy $\pi_{[k]}(\cdot|h(k))$ as $\pi_{[k]}^{h}$. Note that $\pi_{[k]}$ refers to the $k$th policy in the population, and $K$ denotes the policy population size. $\Pi$ can be initialized (Line 1 in Algo. 1) in two ways: lazy initialization and immediate initialization. In lazy initialization, $\Pi$ starts with $K$ placeholder policies, with the actual policies being initialized during the training iterations (Line 4 in Algo. 1). In immediate initialization, $\Pi$ is initialized upfront with $K$ real policies, which may be randomly generated or pre-trained models. A policy $\pi_{[k]}^{h}$ is considered inactive if it is a placeholder; otherwise, it is considered active.
@@
  • $\Sigma:=\{\sigma_{[k]}\}_{k=1}^{K+1}\in\mathbb{R}^{K\times C_{1}}$: The interaction matrix $\Sigma$ consists of rows $\sigma_{[k]}$. Each $\sigma_{[k]}\in\mathbb{R}^{C_{1}}$ represents the probability distribution over opponent policies in the policy population $\Pi$, where $C_{1}=K^{n-1}$ and $n$ denotes the number of players. Specially, in a two-player game, if $C_{1}=K$ and $\Sigma_{mn}$ represents the probability that policy $m$ is optimized against policy $n$, $\Sigma$ can be depicted in directed interaction graphs (Fig. 4). $\Sigma$ will be initialized with all zeros (Line 1 in Algo. 1), providing a starting point before the meta-strategy solver updates it using the first payoff information.
@@
  • For epoch $e\in[[E]]$ (Line 2 in Algo. 1): $E$ denotes the total number of epochs for the entire policy population. For example, if the algorithm only introduces new policies into the population without updating existing ones, then $E=1$. This implies that only the inactive policy is likely to be chosen for training in each iteration and turned into an active one, while active policies remain unchanged. Conversely, if active policies are updated multiple times throughout the algorithm, then $E>1$. Indeed, $E$ reflects the number of updates performed.
@@
  • Initialize $\pi_{[k]}^{h}$ (Line 4 in Algo. 1): The initialization of $\pi_{[k]}^{h}$ can vary depending on the algorithm being used. It can be initialized randomly or by leveraging pre-trained models (Silver et al., 2016; Vinyals et al., 2019) or through a recently updated policy.
@@
  • $\text{ORACLE}(\pi_{[k]},\sigma_{[k]},\Pi)$ (Line 5 in Algo. 1): The ORACLE is an abstract computational entity that returns a new policy adhering to specific criteria. Here, we divide ORACLE into three types. (1) One type is the BR oracle, which is designed to identify the optimal counter-strategies against an opponent’s strategy, including finding NE (McMahan et al., 2003). However, it often requires considerable computational effort. (2) To alleviate the computational demands, the approximate best response (ABR) oracle is introduced, using techniques such as RL (Algo. 2), evolution-theory-based methods (Algo. 3) or regret minimization methods (Algo. 4). (3) Other specially crafted ORACLEs are either tailored to specific meta-strategy solvers (MSSs) (Muller et al., 2020; Marris et al., 2021) or introduced to enhance diversity (Balduzzi et al., 2019; Perez-Nieves et al., 2021; Liu et al., 2021). Some details of Algo. 23 and 4 need to be mentioned. Algo. 2 is a RL method that can be widely used across different categories. Moreover, on-policy RL algorithms do not require a sample buffer to collect trajectories. In contrast, only off-policy methods rely on a sample buffer to store and reuse past experiences (Line 11 and 15 in Algo. 2). Algo. 3 is specifically utilized by Regularized Nash Dynamics (R-NaD) (Perolat et al., 2021) and Algo. 4 is specifically utilized by the regret-minimization-based series. Details are provided in Sec. 3.3.5 and Sec. 3.5, respectively.
@@
  • $\text{EVAL}(\Pi)$ (Line 6 in Algo. 1): Evaluating the policy population $\Pi$. There are multiple evaluation metrics available to assess the performance of each policy. The performance matrix is represented as $\mathcal{P}:=\{p_{k}\}_{k=1}^{K}\in\mathbb{R}^{K\times C_{2}}$. $p_{k}$ is the performance of policy k, and the dimension $C_{2}$ depends on the evaluation metric used. For instance, $p_{k}$ can be depicted as the relative skill like Elo ratings ($C_{2}=1$) or can be depicted as the payoff tensor ($C_{2}=K^{n-1}$, where n is the number of players). Specially, in two-player symmetric zero-sum games, the expected payoffs can serve as the evaluation metric. In such cases, $\mathcal{P}$ is a square matrix ($C_{2}=K$).
@@
  • $\textit{MSS}:\mathbb{R}^{K\times C_{2}}\rightarrow\mathbb{R}^{K\times C_{1}}$ (Line 7 in Algo. 1): A meta-strategy solver (MSS) takes the performance matrix $\mathcal{P}$ as its input and produces a new interaction matrix $\Sigma$ as its output. If the MSS produces a constant matrix irrespective of the input, the evaluation step (Line 6 in Algo. 1) can be skipped to reduce computations. In this case, the MSS step also can be skipped. Additionally, if it is the final iteration, the MSS step should also be skipped, as the resulting meta-strategy will not be used. Moreover, if $\Pi$ uses immediate initialization, $\Sigma$ will be initialized (Line 1 in Algo. 1) by first evaluating the performance and then applying the MSS.
@@
Figure 4. Interaction matrix $\Sigma$ examples. When $C_{1}=K$ and the opponent sampling strategy $\Sigma_{mn}$ represents the probability that policy $m$ is optimized against policy $n$, the interaction matrix $\Sigma$ can be depicted in directed interaction graphs. Here, we consider three representative self-play algorithms. In the Top section, we define the interaction matrix $\Sigma\in\mathbb{R}^{3\times 3}$ as $\{\sigma_{[k]}\}_{k=1}^{3}$. In the Bottom section, we present directed interaction graphs where the outgoing edges from each node are equally weighted, and their weights collectively sum to one. The relationship between the Top and Bottom sections is established through directed edges: an edge directed from node $m$ to node $n$ with a weight of $\Sigma_{mn}$ signifies that policy $m$ is optimized against policy $n$ with a probability of $\Sigma_{mn}$. Note that this figure is reproduced from Liu et al. (2022c) and this concept is initially proposed by Garnelo et al. (2021).
@@
[…]
0:$B$ $\triangleright$ Sample buffer (optional).
1:while $\pi_{[k]}^{h}$ is not valid do
@@
3:   sample $\boldsymbol{\pi_{opp}^{h}}\sim P(\sigma_{[k]})\triangleright$ Policies of opponents.
4:   $\boldsymbol{\pi}=(\pi_{[k]}^{h},\boldsymbol{\pi_{opp}^{h}})$
5:   $s_{0},\boldsymbol{o_{0}}\sim\rho$ $\triangleright$ Initial state.
@@
7:    $\boldsymbol{a_{t}}\sim\boldsymbol{\pi}(\boldsymbol{o_{t}})$
8:    $s_{t+1},\boldsymbol{o_{t+1}}\sim P(s_{t},\boldsymbol{a_{t}})$
9:    $\boldsymbol{r_{t}}\leftarrow\boldsymbol{R}(s_{t},\boldsymbol{a_{t}})$
@@
11:   $B\leftarrow B\cup\{\tau\}$ $\triangleright$ Store the trajectory in the buffer (optional).
12:   $\pi_{[k]}^{h}\leftarrow\text{update}(\pi_{[k]}^{h})\triangleright$ Use RL algorithms to train the policy $k$.
13:  end for
14:end while
15:return $\pi_{[k]}^{h},B$ $\triangleright$ Return the buffer (optional).
@@
[…]
1: sample $\boldsymbol{\pi_{opp}^{h}}\sim P(\sigma_{[k]})\triangleright$ Policies of opponents.
[…]
3: Transform the reward by adding a policy-dependent reward according to $\boldsymbol{\pi_{\text{reg}}}$ (Perolat et al., 2021).
[…]
5:return $\pi_{[k]}^{h}$
@@
[…]
1:for each player i do
2:  sample $\boldsymbol{\pi_{opp}^{h}}\sim P(\sigma_{[k]})\triangleright$ Policies of opponents.
3:  $\boldsymbol{\pi}=(\pi_{[k]}^{h}(i),\boldsymbol{\pi_{opp}^{h}}(-i))$
4:  Use regret matching to play the game and update $h(k)$ with new regret minimization information.
@@
6:return $\pi_{[k]}^{h}$
3.2. Traditional Self-Play Algorithms